# Gravitational waves of quasi-circular, inspiraling black hole binaries in an ultralight vector dark-matter environment

### v1

@@ Appendix A Binary systems in ultralight dark matter halos
In this appendix, we discuss the validity of approximating the dark matter energy profile as homogeneous on the extension of the orbit. Numerical solutions show that dark matter halos in ultralight dark matter models can be described by solitonic cores surrounded by Navarro-Frenk-White profiles. The radial density profile of the core can be modeled as [45, 23, 46]
@@
where $r_{c}$ is the core radius, defined as the radius at which the density drops to half of its peak value. The mass of the core is defined as the mass enclosed by $r_{c}$. For ultralight scalar field models, the core radius can be related to dark-matter field mass and the corresponding total halo masses. At redshift $z=0$, $r_{c}$ can be estimated as [^6] See for instance Eq. (7) in Ref. [45].
$r_{c}=10^{12}\left(\frac{10^{-18}\rm{eV}}{m_{A}}\right)\left(\frac{M_{h}}{10^{9}\rm{M}_{\odot}}\right)^{-1/3}\rm{km}\,.$ (64)
If we suppose that each halo hosts a galaxy with a supermassive black at its center [47, 48], we can further relate the mass of the black hole with the mass of the dark matter halo as [49, 50, 51]
@@
In order to explain our set up, we consider the case of supermassive and low/intermediate massive binaries separately. In the case of supermassive black hole binaries, it is reasonable to assume that, due to processes such as dynamical friction and mass segregation, the black holes are most probably located near the center of their host galaxy. Moreover, these galaxies themselves are typically located near the core centers of their respective halos [52, 53], where the energy density is approximately constant (see Eq. 62). Then, if the size of the orbit $b$ is much smaller than the size of the core $r_{c}$, we can assume that the energy profile throughout the binary is approximately homogeneous. To check the validity of this assumption, we consider a supermassive binary of $M=10^{6}\rm{M}_{\odot}$, which corresponds to the largest system LISA can probe. Furthermore, using Eq. (65) for this system with a field environment of mass $m_{A}=10^{-18}\rm{eV}$, we obtain a halo mass of approximately $M_{\rm{halo}}=10^{11}\rm{M}_{\odot}$, with a corresponding central core of radius $r_{c}=2.5\times 10^{11}\rm{km}$. We thus see that the core size is much larger than the orbits considered in this work ($b\sim 10^{9}\rm{km}$), so we can assume an approximately homogeneous energy density profile in the extension of the orbit. Also, we see that this estimate matches the one from the discussion in Sec. II regarding the de Broglie wavelength given in Eq. (5), since we used the velocity dispersion of the Milky Way, which hosts a supermassive black hole at its center (Sag. ${\rm{A}}^{\star}$) of $\sim 4\times 10^{6}\rm{M}_{\odot}$, and has a halo mass of $\sim 7\times 10^{11}\rm{M}_{\odot}$.
For supermassive binaries, there is also non-negligible backreaction of the binary on the dark-matter field environment. Numerical simulations of inspiralling supermassive binaries in ultralight dark-matter field environments show that this backreaction generates a quasi-stationary energy-density profile around the binary, which is approximately homogeneous throughout the extension of the orbit, and is peaked in the vicinity of the black holes (see for instance Refs. [54, 55, 56]). This energy profile can be understood as a local enhancement of the host soliton density profile, produced by the accretion of dark matter by the black holes. We here ignore this effect, so the values used for the density of the dark matter field in the orbit are in general underestimated. A larger density would only enhance the effects considered here, making our estimates necessarily conservative.

### v2

@@ Appendix A Binary systems in ultralight dark matter halos
In this appendix, we discuss the validity of approximating the dark matter energy profile as homogeneous on the extension of the orbit. Numerical solutions show that dark matter halos in ultralight dark matter models can be described by solitonic cores surrounded by Navarro-Frenk-White profiles. The radial density profile of the core can be modeled as [Schive:2014hza, Schive:2014dra, Schwabe:2016rze]
@@
where $r_{c}$ is the core radius, defined as the radius at which the density drops to half of its peak value. The mass of the core is defined as the mass enclosed by $r_{c}$. For ultralight scalar field models, the core radius can be related to dark-matter field mass and the corresponding total halo masses. At redshift $z=0$, $r_{c}$ can be estimated as [^7] See for instance Eq. (7) in Ref. [Schive:2014hza].
$r_{c}=10^{11}\left(\frac{10^{-18}\rm{eV}}{m_{A}}\right)\left(\frac{M_{\rm{halo}}}{10^{12}\rm{M}_{\odot}}\right)^{-1/3}\rm{km}\,.$ (67)
If we suppose that each halo hosts a galaxy with a supermassive black at its center [Kormendy:1995er, Ferrarese:2000se], we can further relate the mass of the black hole with the mass of the dark matter halo as [Ferrarese:2002ct, 10.1111/j.1745-3933.2010.00832.x, Bandara_2009]
@@
In order to explain our set up, we consider the case of supermassive and low/intermediate massive binaries separately. In the case of supermassive black hole binaries, it is reasonable to assume that, due to processes such as dynamical friction and mass segregation, the black holes are most probably located near the center of their host galaxy. Moreover, these galaxies themselves are typically located near the core centers of their respective halos [Kravtsov:2003sg, Berlind:2002rn], where the energy density is approximately constant (see Eq. 65). Then, if the size of the orbit $b$ is much smaller than the size of the core $r_{c}$, we can assume that the energy profile throughout the binary is approximately homogeneous. To check the validity of this assumption, we consider a supermassive binary of $M=10^{6}\rm{M}_{\odot}$, which corresponds to the largest system LISA can probe. Using Eq. (68) for this system we obtain a halo mass of approximately $M_{\rm{halo}}=10^{11}\rm{M}_{\odot}$; and from Eq. (67), for a field environment of mass $m_{A}=10^{-18}\rm{eV}$, we obtain a corresponding central core of radius $r_{c}=2.5\times 10^{11}\rm{km}$. We thus see that the core size is much larger than the orbits considered in this work (see Eq. 6), so we can assume an approximately homogeneous energy density profile in the extension of the orbit. Also, we see that this estimate matches the one from the discussion in Sec. II regarding the de Broglie wavelength given in Eq. (5), since we used the velocity dispersion of the Milky Way, which hosts a supermassive black hole at its center (Sag. $\rm{A}$$^{\star}$) of $\sim 4\times 10^{6}\rm{M}_{\odot}$, and has a halo mass of $\sim 7\times 10^{11}\rm{M}_{\odot}$.
For supermassive binaries, there is also non-negligible backreaction of the binary on the dark-matter field environment. Numerical simulations of inspiralling supermassive binaries in ultralight dark-matter field environments show that this backreaction generates a quasi-stationary energy-density profile around the binary, which is approximately homogeneous throughout the extension of the orbit, and is peaked in the vicinity of the black holes (see for instance Refs. [Bamber:2022pbs, Aurrekoetxea:2023jwk, Boey:2025qbo]). This energy profile can be understood as a local enhancement of the host soliton density profile, produced by the accretion of dark matter by the black holes. The effect is maximal when the orbital radius $r_{12}$ is of order of the Comptom wavelenth $\lambda_{c}$ of the field. In our case, the orbital radius of the systems satisfies $r_{12}\ll\lambda_{c}$. We here ignore this effect, so the values used for the density of the dark matter field in the orbit are in general underestimated. A larger density would only enhance the effects considered here, making our estimates necessarily conservative.